a photon of wavelength 300 nm is absorbed by a gas and then re-emitted as two photons, one photon is red with wavelength 760 nm. The wave number of second photon will be

Grade:12th pass

1 Answers

Aman gohel
6 years ago
wavelength of absorbed photon=300nm,E=hc/lamda,According to law of conservation of energy,=>E=E1+E2,=>(hc/300)=(hc/760)+hc/lamda2hc got cancelled,1/300=1/760+1/lamda2=>1/lamda2=23/11400=Therefore lamda2=11400/23=[495.65nm](Ans)